find the enthalpy of entropy of steam when the pressure is 20 Bar and specific volume is 0.8
1
Expert's answer
2020-12-22T05:07:44-0500

The Dryness fraction(x) of the wet steam=0.8


The pressure of wet steam(P)=20 bar=2 MPa J/kg


The saturation properties of wet steam at 2 MPa is given


in the steam table as

"V= 0.0012 m3\/kg Vtg (0.0996-0.0012)=0.0984 m3\/kg"

"h=908.5 kJ\/kg hig=1889.8kJ\/kg"

"U = 906.15 kJ\/kg Ufq= (2599.1-906.15) =1692.95 kJ\/Kg S= 2.4468 kJ\/kg-K s=3.8923 KJ\/kg-K"

Specific volume at "x=0.8 = v=vf+x.vfg=0.07992 m3\/kg"

Specific enthalpy at "x=0.8 => h=hf+x.hfg=2420.34 kJ\/kg"

Specific internal energy at "x=0.8=>u=uf+x.ufg=2260.51 k"

Specific entropy at "x=0.8 => s=sf+x sfg=5 5606 kJ\/kg-K"